Growth​ Company's current share price is $20.25 and it is expected to pay a $0.95 dividend per share next year. After​ that, the​ firm's dividends are expected to grow at a rate of 4.4% per year.

a. What is an estimate of Growth​ Company's cost of​ equity?

b. Growth Company also has preferred stock outstanding that pays a $2.15 per share fixed dividend. If this stock is currently priced at $28.15​, what is Growth​ Company's cost of preferred​ stock?c. Growth Company has existing debt issued three years ago with a coupon rate of 5.8%.

The firm just issued new debt at par with a coupon rate of 6.2%. What is Growth​ Company's cost of​ debt?d. Growth Company has 5.3 million common shares outstanding and 1.2 million preferred shares​ outstanding, and its equity has a total book value of $50.3 million. Its liabilities have a market value of $20.2 million. If Growth​ Company's common and preferred shares are priced as in parts ​(a​) and ​(b​), what is the market value of Growth​ Company's assets?e. Growth Company faces a 38% tax rate. Given the information in parts ​(a​) through ​(d​), and your answers to those​ problems, what is Growth​ Company's WACC?

​Note: Assume that the firm will always be able to utilize its full interest tax shield.

a. What is an estimate of Growth​ Company's cost of​ equity?

The required return​ (cost of​ capital) of levered equity is

nothing​%.

​(Round to two decimal​ places.)b. Growth Company also has preferred stock outstanding that pays a $2.15 per share fixed dividend. If this stock is currently priced at $28.15​, what is Growth​ Company's cost of preferred​ stock?The cost of capital for preferred stock is

nothing​%.

​(Round to two decimal​ places.)c. Growth Company has existing debt issued three years ago with a coupon rate of 5.8%. The firm just issued new debt at par with a coupon rate of 6.2%.

What is Growth​ Company's cost of​ debt? ​ (Select from the​ drop-down menus.)The​ pre-tax cost of debt is the​ firm's YTM on current debt. Since the firm recently issued debt at​ par, then the coupon rate of that debt must be

less than

equal to

greater than

the YTM of the debt.​ Thus, the​ pre-tax cost of debt is

6.2%

5.8%

. d. Growth Company has 5.3 million common shares outstanding and 1.2 million preferred shares​ outstanding, and its equity has a total book value of $50.3 million. Its liabilities have a market value of $20.2 million. If Growth​ Company's common and preferred shares are priced as in parts ​(a​) and ​(b​), what is the market value of Growth​ Company's assets?The market value of assets is

​$nothing

million.  ​(Round to two decimal​ places.)e. Growth Company faces a 38% tax rate. Given the information in parts ​(a​) through ​(d​), and your answers to those​ problems, what is Growth​ Company's WACC?The weighted average cost of capital is

nothing​%.

​(Round to two decimal​ places.)
Answer #1

a) Cost of equity, re = D1 / P0 + g = 0.95 / 20.25 + 4.4% = 9.09%

b) Cost of preferred, rp = D / P = 2.15 / 28.15 = 7.64%

c) Cost of debt, rd = 6.2%

d)

Value Weight Cost
Equity 107.3 66.5% 9.09%
Preferred 33.8 20.9% 7.64%
Debt 20.2 12.5% 6.20%
Total 161.3 WACC 8.13%

Value = no. of shares x price

Weight = Value / Total

WACC = wd x rd x (1 - tax) + wp x rp + we x re

= 8.13%